OpenStudy (ashleyisakitty):

Please help me! AB is tangent to circle O at B. Find the length of the radius r for AB=8 and AO=11.7.

OpenStudy (ashleyisakitty):

AB^2=BO^2-AO^2 =(BC+CO)^2-AO^2 =(11.7+8)^2-(8)^2 (19)^2-(8)^2 (19+8)(19-8) (27)(11) sqrt(27*11)=sqrt(27)*sqrt(11)=5.1*3.3= =16.8 i did this but my answer doesnt fit any of my choices

OpenStudy (anonymous):

10 It's a right triangle, a^2+36=136.89 so 10.04

OpenStudy (amistre64):